A diverging mirror produces what kind of images? None; they cannot produce images Both real and virtual images Only real images Only virtual images Submit Request Answer
Answer #1

Here ,

for a diverging mirror

as the focal length is positive

and 1/f =1/v + 1/u

1/v = 1/f -1/u

as u is negative

1/v is positive and , v is positive

hence , the image is always virtual

Only virtual images